OpenStudy (anonymous):

i dont understand very well how to do this Choose one of the factors of x6 + 1000 x2 x2 – 10 x4 – 10x2 + 100 x4 + 10x2 + 100

OpenStudy (jdoe0001):

\(\bf x^6+1000\) has 2 factors, one of those listed is one of them

OpenStudy (jdoe0001):

keep in mind that \(\bf x^6 = (x^2)^3\\ 1000 = 10^3\\ \color{blue}{a^3+b^3 = (a+b)(a^2-ab+b^2)}\)

OpenStudy (anonymous):

what does the "bf" and the color blue mean?

OpenStudy (jdoe0001):

hehe, means you're using a rather older browser or internet explorer I think

OpenStudy (jdoe0001):

try this same page in another browser

OpenStudy (anonymous):

i dont have another browser just internet explorer :/

OpenStudy (jdoe0001):

ughh!

OpenStudy (anonymous):

i tried getting google chrome but it always failed to load

OpenStudy (jdoe0001):

that's really terrible, IE is not so great sadly

OpenStudy (jdoe0001):

download something like google chrome

OpenStudy (jdoe0001):

or firefox

OpenStudy (anonymous):

is there another way to explain it?

OpenStudy (jdoe0001):

any good reason on not getting a better browser? other than loyalty maybe

OpenStudy (anonymous):

i have no other broser

OpenStudy (jdoe0001):

I could write just like say x^6 = (x^2)^3 10 = 10^3 (a^3+b^3)=(a+b)(a^2−ab+b^2) I guess

OpenStudy (anonymous):

that works

OpenStudy (jdoe0001):

well, google chrome doesn't cost anything nor does firefox, sounds to me you have no good reasons not to get either

OpenStudy (jdoe0001):

needless to say they're far better than IE anyway not trying to knock down IE, but I do know the standards, IE lags behind several major browsers

OpenStudy (mathmate):

Many times Google chrome is unable to load openstudy when the forum is up and running...

OpenStudy (mathmate):

The same goes for an old IE.

OpenStudy (mathmate):

I have firefox and the other two running to see which one actually works.

OpenStudy (jdoe0001):

well, that's is true, but it has nothing to do with the browser as far as I know is just the webserver at openstudy that has a few latency issues

OpenStudy (jdoe0001):

I know this because I've test it in 4 browsers, same site, same connection

OpenStudy (jdoe0001):

when you have a page having a hard time loading in your browser, try your connection on another server, you'll notice the other servers load quick, whilst openstudy is lagging some they've improved a bit btw, it used to worse

OpenStudy (anonymous):

i have google chrome now

OpenStudy (mathmate):

Congrats! If your computer and Internet connection are both fast enough, you could fire up 2 or three browsers and have more chances of getting it to work, if they connect to different servers, as @jdoe0001 mentioned.

OpenStudy (jdoe0001):

so this is what I typed earlier \(\bf x^6 = (x^2)^3\\ 1000 = 10^3\\ \color{blue}{a^3+b^3 = (a+b)(a^2-ab+b^2)}\)

OpenStudy (anonymous):

yeah i see it now

OpenStudy (jdoe0001):

so, do you see the factor then?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

ok